Bibliography file Moon.bib

Version of Wed 17-Mar-93 13:55

March 17, 1993

Bibliography

[1] J. H. Allton and H.V. Lauer. Robotic geology field trip on mars: The preservation of martian rock and soil samples. In S. Johnson and J. Wetzel, editors, Engineering, Construction, and Operations in Space II. ASCE International Conference, Albequerque NM, April 1990. Full index of proceedings available.

Key: ALLTON90A

[2] Victor R. Baker, Ronald Greeley, Paul D. Komar, Donald A. Swanson, and Richard B. Waitt Jr. Columbia and snake river plains. In Geological Society of America Centennial Special Volume 2, chapter 11. Geological Society of America, 1987.

Key: BAKER87A

[3] Leonhard E. Bernold and Shankar Sundareswaran. A study on the mechanics of lunar excavation. Technical report, Department of Civil Engineering, University of Maryland College Park, MD 20742.

Key: BERNOLDA

[4] John Carmody and Raymond L. Sterling. Underground space on earth: Analog for a lunar base. In S. Johnson and J. Wetzel, editors, Engineering, Construction, and Operations in Space II. ASCE International Conference, Albequerque NM, April 1990. Full index of proceedings available.

Key: CARMODY90A

[5] Cassandra R. Coombs and B. Ray Hawke. A search for intact lava tubes on the moon: Possible lunar base habitats. In W. W. Mendell, editor, The Second Conference on Lunar Bases and Space Activities of the 21st Century, April 1988.

Key: COOMBS88B

[6] Cassandra R. Coombs, B. Ray Hawke, and Lionel Wilson. Geologic and remote sensing studies of rima mozart. In Proceedings of the Eighteenth Lunar and Planetary Science Conference. Cambridge University Press, 1988.

Key: COOMBS88A

[7] Andrew W. Daga, Meryl A. Daga, and Wendel R. Wendel. A preliminary assessment of the potential of lava tube-situated lunar base architecture. In S. Johnson and J. Wetzel, editors, Engineering, Construction, and Operations in Space II. ASCE International Conference, Albequerque NM, April 1990. Full index of proceedings available.

Key: DAGA90A

[8] Joseph Fieber. Lunar structures. Lunar Toons, 2(2), 1991.

Key: FIEBER91A

[9] Ronald Greeley. Lava tubes and channels in the lunar marius hills. The Moon, An international journal of Lunar studies, May 1971.

Key: GREELEY71A

[10] Ronald Greeley. Lunar hadley rille: Considerations of its origin. Science, 172:722–725, May 1971.

Key: GREELEY71B [11] Ronald Greeley. Modes of emplacement of basalt terrains and an

analysis of mare volcanism in the orientale basin. In Proceedings Lunar Science Conf. 7th, pages 2747–2759, 1976.

Key: GREELEY72E

[12] Ronald Greeley. The snake river plain, idaho: Representative of a new category of volcanism. Journal of Geophysical Research, 87(B4):2705–2712, April 1982.

Key: GREELEY82H

[13] Ronald Greeley. The style of basaltic volcanism in the eastern snake river plain, idaho. Cenozoic Geology of Idaho: Idaho Bureau of Mines and Geology Bulletin, (26):407–421, 1982.

Key: GREELEY82I

[14] Ronald Greeley. The role of lava tubes in hawaiian volcanoes. In Volcanism in Hawaii, chapter 59. Hawaiian Volcano Observatory, 1987.

Key: GREELEY87J

[15] Ronald Greeley and D. Gault. Observations of actively forming lava tubes and associated structures, hawaii. Modern Geology, 2:207–223, 1971.

Key: GREELEY71C

[16] Ronald Greeley and D. Gault. Additional observations of actively forming lava tubes and associated structures, hawaii. Modern Geology, 3:157–160, 1972.

Key: GREELEY72D

[17] Ronald Greeley and Donald E. Gault. Endogenic craters on basaltic lava flows: Size frequency distributions. In Proceedings Lunar Planet Sci.ence Conf. 10th, pages 2919–2933, 1979.

Key: GREELEY79G [18] Ronald Greeley and Paul D. Spudis. Mare volcanism in the herigo-

nius region of the moon. In Proceedings Lunar Planet Sci.ence Conf. 9th, pages 3333–3349, 1978.

Key: GREELEY78F

[19] John E. Gruener. El malpais lava tubes: Analog to lunar lava tubes. In S. Johnson and J. Wetzel, editors, Engineering, Construction, and Operations in Space II. ASCE International Conference, Albequerque NM, April 1990. Full index of proceedings available.

Key: GRUENER90A

[20] J.E. Guest, C. Wood, and R. Greeley. Lava tubes, terraces and megatumuli on the 1614-24 pahoehoe lava flow field, mount etna, sicily. Bulletin Volcanologique, 47(3), 1984.

Key: GUEST84A

[21] Friedrich Horz. Lava tubes: Potential shelters for habitats. In W. W. Mendell, editor, Lunar Bases and Space Activities of the 21st Century, pages 405–412. Lunar and Planetary Institute, Houston. Full index of the proceedings available.

Key: HORZA

[22] W. David Carrier III and James K. Mitchell. Geotechnical engineering on the moon.

Key: CARRIERA

[23] E. Nader Khalili. Lunar, martian and asteroid structures built in situ. In S. Johnson and J. Wetzel, editors, Engineering, Construction, and Operations in Space II. ASCE International Conference, Albequerque NM, April 1990. Full index of proceedings available.

Key: KHALILI90A [24] Paul R. Klarer. Extending terrestrial mobile robotics to a pro-

posed lunar base. In S. Johnson and J. Wetzel, editors, Engineering, Construction, and Operations in Space II. ASCE International Conference, Albequerque NM, April 1990. Full index of proceedings available.

Key: KLARER90A

[25] Lunar, Mars Exploration Office, and Planet Surface Systems Office. Planet surface systems study period summary. Technical report, NASA, November 1989.

Key: NASA89A

[26] John S. Rinehart and Ronald Greeley. Seismic-wave velocity patterns in some pahoehoe basalt flows. Journal of Geophysical Research, 76(23), August 1971.

Key: RINEHART71A

[27] Peter H. Schultz, Ronald Greeley, and Donald E. Gault. Degradation of small mare surface features. In Proceedings Lunar Science Conf. 7th, pages 985–1003, 1976.

Key: SCHULTZ76A

[28] Paul D. Spudis, Gordan A. Swann, and Ronald Greeley. The formulation of hadley rille and implications for the geology of the apollo 15 region. In Proceedings of the eighteenth lunar and planetary science conference, 1988.

Key: SPUDIS88B

[29] Paul D. Spudis and G. Jeffrey Taylor. Rationale and requirements for lunar exploration. In S. Johnson and J. Wetzel, editors, Engineering, Construction, and Operations in Space II. ASCE International Conference, Albequerque NM, April 1990. Full index of proceedings available. Spudis : 713-486-2192.

Key: SPUDIS90A [30] G. Jeffrey Taylor and Paul D. Spudis. A teleoperated robotic field

geologist. In S. Johnson and J. Wetzel, editors, Engineering, Construction, and Operations in Space II. ASCE International Conference, Albequerque NM, April 1990. Full index of proceedings available.

Key: TAYLOR90A

[31] E. Theilig and R. Greeley. Lava flows on mars: Analysis of small surface features and comparisons with terrestrial analogs. In Proceedings of the seventeenth lunar and planetary science conference, Part 1 Journal of geophysical research, volume 91, pages E193– E206, November 1986.

Key: THEILIG86A

[32] L. D. Toups. Reference architecture description for the lunar/mars 90 day study period. Nasa contract nas9-17900, Lockheed Engineering and Sciences Company, Inc., January 1990.

Key: TOUPS90A

[33] Larry Toups. A survey of lunar construction techniques. In S. Johnson and J. Wetzel, editors, Engineering, Construction, and Operations in Space II. ASCE International Conference, Albequerque NM, April 1990. Full index of proceedings available.

Key: TOUPS90B